OpenStudy (anonymous):

Complete the following ordered pairs for the equation. 6x + -5y = 20 (-5, _) (_, -4) (_, 2)

OpenStudy (hunus):

Ordered pairs are of the form (x, y) So try plugging in the value your given in an ordered pair into its corresponding variable and work out what the other one should be

OpenStudy (anonymous):

i need answers :( because i am dumb

OpenStudy (hunus):

You aren't dumb. I'll do the first one for you so you know how to solve them. You're given x = -5 So plug -5 into x in the equation you're given 6(-5) + -5y = 20 -30 - 5y = 20 -5y = 50 y = -10 So that ordered pair would be(-5, -10)

OpenStudy (anonymous):

thankyou i understand :)

OpenStudy (hunus):

Good :)

OpenStudy (anonymous):

wait how didi you did it on the last step -5y=50 how did you got -10?

OpenStudy (hunus):

a negative times a negative is a positive and the question is -5 times what is 50? So -5 * 10 would be - 50 but we need positive 50 so it's -5 * -10